PA°equal to 200 torr and PB°equals to 100 torr calculate the composition of vapour of liquid solution containing A and B at 50 degree centigrade is the normal boiling point of solution?
1
Expert's answer
2020-05-19T08:43:39-0400

Solution.

"P = P^0(A) + P^0(B)"

P = 300 torr

"\chi(A) = \frac{P^0(A)}{P} = 0.67"

"\chi(B) = \frac{P^0(B)}{P} =0.33"

"t = \chi(a) \times P^0(A) + \chi(B) \times P^0(B)"

"t = 0.67 \times 200 + 0.33 \times 100 = 167 \ K"

50 is not a normal boiling point.

Answer:

"\chi(A) = \frac{P^0(A)}{P} = 0.67"

"\chi(B) = \frac{P^0(B)}{P} =0.33"

50 is not a normal boiling point.